OpenStudy (anonymous):

A principal of $700 is invested in an account at 6% per year compounded annually. What is the total amount of money in the account after 5 years? A. $910.00 B. $920.00 C. $936.76 D. $928.84

OpenStudy (solomonzelman):

\(\LARGE\color{slate}{ \displaystyle {\rm F} = {\rm P} \left( 1+\frac{r}{n}\right) ^{n \times t} }\) ``` F is the future value. P is the principal/current amount. n is the number of times the interest is compound per year. r is the percent rate (but, in decimal, so if you had 3%, then you would write 0.03 ) t is the number of years of the investment. ```

OpenStudy (anonymous):

Ok when i put this into mathway.com maybe i put it in wrong but i do not get an answer even close to the answers it gives me.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\LARGE\color{slate}{ \displaystyle {\rm F} = {\rm P} \left( 1+\frac{r}{n}\right) ^{n \times t} }\) \(\LARGE\color{slate}{ \displaystyle {\rm F} = {\rm 900} \left( 1+\frac{0.06}{1}\right) ^{1 \times 5} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{slate}{ \displaystyle {\rm F} = {\rm 900} \left( 1.06\right) ^{ 5} }\)

OpenStudy (anonymous):

Ohh

OpenStudy (solomonzelman):

oh, my bad

OpenStudy (solomonzelman):

everything is correct, but instead of 900, it is 700

OpenStudy (solomonzelman):

because you are investing 700 in the question, and not 900.

OpenStudy (solomonzelman):

confused, or good ?

OpenStudy (solomonzelman):

\(\LARGE\color{slate}{ \displaystyle {\rm F} = {\rm 700} \left( 1.06\right) ^{ 5} }\)

OpenStudy (anonymous):

I think i got it.. it would be C

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

very good!